OpenStudy (anonymous):

Please help. I will give medal to best answer. Choose the equation below that represents the line passing through the point (-5, 1) with a slope of three halves. y - 5 = three halves(x + 1) y + 1 = three halves(x - 5) y + 5 = three halves(x - 1) y - 1 = three halves(x + 5)

OpenStudy (anonymous):

Can you please help me @binks?

OpenStudy (anonymous):

PLEASE HELP me!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (zpupster):

slope-intercept equation of all lines is y=mx+b A point-slope equation of lines is y-y1 = m(x-x1) m=slope (x1,y1) is the ordered pair just plug them in and watch your signs

OpenStudy (zpupster):

what did you get?

OpenStudy (anonymous):

I'm sorry that just totally confused me.

OpenStudy (zpupster):

slow down and breathe i will show you how to plug them in- (-5,1) = (x1,y1) and m=3/2 here is our equation y-y1 = m(x-x1) y-(-5) = 3/2(x-(1) now what did you get??

OpenStudy (anonymous):

I don't know how to solve it but my answer is the third on which is y + 5 = 3/2(x-1) right?

OpenStudy (zpupster):

good job!!

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

sorry just got this glad you got it